The equation of a straight line which bisects the intercepts made by the axes on the lines x+y=2 and 2x+3y=6 is?

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given lines arelines x+y=2 and 2x+3y=6
so the intercepts are (2,0) and (0,2)
and (3,0) and (0,2)
SoThe equation of a straight line which bisects the intercepts made by the axes on the lines x+y=2 and 2x+3y=6 is given by y=3-2 which is y=1